1. If A + B = 45° then (1+ tanA)(1+ tanB)=
2. Let A and B be such that π
 

Grade:11

1 Answers

Arun
5 years ago
Dear Ayushi
 
tan (A + B) = tan 45 = tan A + tanB / 1 – tanA tanB
 
1 – tan A tanB = tan A + tanB
 
1 + 1 = 1 + tan A  + tan B + tanA tanB
 
2 = (1 + tanA) (1 + tanB)
 
Regards
Arun